OpenStudy (ladylauren):

A pair of equations is shown below: y = 7x − 9 y = 3x − 1 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations. (6 points) Part B: What is the solution to the pair of equations? (4 points)

OpenStudy (texmechanica):

(2,5)

OpenStudy (ladylauren):

I think the answer to part B is (2,5) but i need to know how to get there so i can explain it.

OpenStudy (texmechanica):

Yes siree.

OpenStudy (texmechanica):

y=mx+b m=slope b= y intercept y=7x-9 slope=7 yintercept=-9 y=3x-1 slope=3 yinercept=-1

OpenStudy (texmechanica):

Then simplify and subtract them to solve.

OpenStudy (texmechanica):

7x-9=y=3x-1 7x-9=3x-1 4x-9=-1 4x=8 x=2 y=3x-1 y=3*2-1 y=6-1 y=5

OpenStudy (texmechanica):

It would be (2,5)

OpenStudy (ladylauren):

ok how would you explain that?

OpenStudy (ladylauren):

because part A is In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations.

OpenStudy (texmechanica):

Take what I said, and then try to put those in your own words.

OpenStudy (texmechanica):

My own words won't be your own words.

OpenStudy (ladylauren):

i just want to know what part it wants like what point or how i got that point.